calculate the two longest wavelength of the raditions emitted when hydrogen atoms makes transitions from higher state to n=2 state

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

The two longest wavelengths would be for transitions corresponding to least energy. Hence the two wavelengths correspond to transitions3\rightarrow 2and4\rightarrow 2.

Hence\dfrac{hc}{\lambda_1}=13.6eV(\dfrac{1}{2^2}-\dfrac{1}{3^2})=1.88eV

and\dfrac{hc}{\lambda_2}=13.6eV(\dfrac{1}{2^2}-\dfrac{1}{4^2})=2.55eV

\implies \lambda_1=659.5nm,\implies \lambda_2=486.3nm